What is the molarity of a solution if 47.50 g potassium iodide (KI) is added to a container that is then filled with water to a final volume of 201 mL?


1
Expert's answer
2021-11-30T01:46:02-0500

Molar mass of KI = 166

= 47.5/166=0.29mol

= 0.29/0.201=1.443mol/L